OpenStudy (anonymous):

can anyone help me that is in algebra 1?

OpenStudy (pinksapphire):

What's the question?

OpenStudy (kohai):

I can help :) What is your question?

geerky42 (geerky42):

@ashleychase1

OpenStudy (anonymous):

do you guys take algebra 1 or not? @PinkSapphire @kohai @geerky42

geerky42 (geerky42):

Last class I have taken is multivariable calculus, of curse I can do algebra :)

OpenStudy (anonymous):

simplify the radical expression √7 + 3 √7

OpenStudy (anonymous):

simplify the radical expression 2√6 + 3√96

OpenStudy (anonymous):

simplify the radical expression (8+ √11)(8- √11)

geerky42 (geerky42):

First one, what do both terms have in common?

OpenStudy (anonymous):

i dont know how to do these at all

geerky42 (geerky42):

Both terms have something you can factor out.

geerky42 (geerky42):

\[\sqrt{7} + 3\sqrt7 \rightarrow 1\sqrt7+3\sqrt7\rightarrow~?\]

OpenStudy (anonymous):

2 √14?

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Familiar with distributive property? \(a(b+c) = ab+ac\)

OpenStudy (pinksapphire):

Think of it like...1x+3x

OpenStudy (pinksapphire):

\[1\sqrt{7}+3\sqrt{7}=4\sqrt{7}\]